OpenStudy (blondie_b23):

Please help me find the distance between -1/6 and 17/12 on a number line.

OpenStudy (blondie_b23):

@zepdrix

OpenStudy (blondie_b23):

@zepdrix

zepdrix (zepdrix):

Do you understand how to rewrite -1/6 in terms of twelfths? :) We would like a common denominator between the two values, a 12 seems like a good choice.

zepdrix (zepdrix):

\[\large\rm -\frac{1}{6}=-\frac{?}{12}\]

OpenStudy (triciaal):

is that 17/12 or 1 and 7/12?

OpenStudy (blondie_b23):

its is 7/12

OpenStudy (blondie_b23):

I'm so confused!!

OpenStudy (mathmale):

First, some practice. Draw a number line. Plot x=-1 and x=13 on this line. What is the distance between these two x-values? Show all work you do.

OpenStudy (blondie_b23):

I really don't understand what this is asking.

OpenStudy (mathmale):

Have you any notes from class about this type of math problem? do you have a textbook, and, if so, have you read the relevant section? Sorry to put you on the spot, but you continue to state "i don't understand." Rather than expect someone else to start from ground zero in explaining the concept of distance between two points on a number line, could you possibly ask questions...taking some control of your own learning?

OpenStudy (mathmale):

In other words, share whatever you have done so far to help yourself learn this material.

OpenStudy (blondie_b23):

You know what? I have spent a lot of time on this site and normally people are polite and helpful. My teacher genuinely didn't explain this topic in class. I understand how it would work in a normal situation. I thought when you were giving the example above, that you were explaining how to do this problem. I didn't understand where you got 1 and 13. I simply don't know how to do this type of question when decimals are involved.

OpenStudy (blondie_b23):

Furthermore, to answer your question, I CAME here to get help and further my learning on the topic. This is a question on a NON-CREDITED ASSIGNMENT!

OpenStudy (mathmale):

I'm sorry if you don't feel I'm "polite" or "helpful." I was, as you guessed, giving you an example problem to work on. I asked you to draw a number line and then plot two points on it: -1 and 13. Could you do that, please?

OpenStudy (mathmale):



OpenStudy (blondie_b23):

The distance is 16

OpenStudy (mathmale):

The distance from x=-1 to x=0 is actually 1; that from 0 to 13 is 13. The sum of these two distances is what?

OpenStudy (mathmale):

Remember, you are measuring the straight line distance from x=-1 to x=13.

OpenStudy (mathmale):

x=0 happens to be between x=-1 and x=13.

OpenStudy (mathmale):



OpenStudy (mathmale):

You could also do this problem by writing the following: 13-(-1). What's that equal to?

OpenStudy (blondie_b23):

14?

OpenStudy (mathmale):

Yes, you've got it. Any questions so far?

OpenStudy (blondie_b23):

No

OpenStudy (mathmale):

Now draw another number line and mark -3/2 and +7/2 on it. Don't worry about where I got that -3/2 or that 7.2; this is strictly an example.

OpenStudy (blondie_b23):

So, can I solve just by doing 17/12 +1/6?

OpenStudy (blondie_b23):

Sorry. I just think I have it now.

OpenStudy (mathmale):

Assuming that you know how to find the LCD, yes, can solve the given problem just by adding those fractions together.

OpenStudy (mathmale):

Do it, please, sharing your work along the way.

zepdrix (zepdrix):

Generally speaking, to find the distance between two numbers, `you subtract the values`. But yes, since you're subtracting a negative, the operation changes to addition. :)\[\rm \text{distance between }-\frac16\text{ and }\frac7{12}\quad=\quad \frac7{12}-\left(-\frac{1}{6}\right)\]

OpenStudy (mathmale):

As zepdrix suggested earlier, use 12 as your LCD. What is 17/12 - (-2/12)?

OpenStudy (mathmale):

Taking this problem one further step towards solution, find the value of \[\frac{ 17-(-2) }{ 12 }\]

OpenStudy (mathmale):

You can combine those 2 fractions because they share the denominator, 12.

OpenStudy (mathmale):

Come on, show your stuff. I have to get off the 'Net now, but will be back on OpenStudy later on.

OpenStudy (blondie_b23):

I figured it out. Thanks. Bye.